# On a combinatorial classification of fine compactified universal Jacobians

### v1

@@ 2.2. Universal and c-stability conditions
By Proposition 2.17, $\mathfrak{m}_{(e,h,A)}^{\Gamma}=\mathfrak{m}_{(e,h,A)}^{\Gamma}(\phi)$ if and only if
(2.7) $\Bigl{\lceil}\frac{(e+2h-2)(d+1-g)+(2g-e-2h)\phi_{A}-(e+2h-2)\phi_{A^{\mathsf{c}}}}{2g-2}\Bigr{\rceil}+h-1=\mathfrak{m}_{(e,h,A)}^{\Gamma},$
while $\mathfrak{m}^{\Gamma}_{(e,g+1-e-h,A^{\mathsf{c}})}=\mathfrak{m}^{\Gamma}_{(e,g+1-e-h,A^{\mathsf{c}})}(\phi)$ if and only if
(2.8) $\Bigl{\lceil}\frac{(-e+2g-2h)(d+1-g)+(e+2h-2)\phi_{A^{\mathsf{c}}}-(2g-e-2h)\phi_{A}}{2g-2}\Bigr{\rceil}+g-e-h=\mathfrak{m}^{\Gamma}_{(e,g+1-e-h,A^{\mathsf{c}})}.$
@@
Clearly, if we choose a family $\{\phi^{\Gamma}\}_{\Gamma\in G_{g,n}}$ defining a universal classical stability condition, the same proof applied to all $(e,h,A)\in\mathcal{D}_{g,n}$ shows that we can check if a universal stability condition is classical on half the vertices of the vine graphs.
@@ 3.1. Curves and dual graphs
We recall the curve-theoretic notation we will use throughout the paper. We refer to [ACGH13] and [PT23] for a detailed exposition.
@@
A curve over an extension $K$ of $k$ is a $\textup{Spec}(K)$-scheme $X$ proper over $K$, geometrically connected and of pure dimension $1$. A curve $X$ is nodal if it is geometrically reduced and, when passing to an algebraic closure $\overline{K}$, its local ring at every singular point is isomorphic to $\overline{K}[[x,y]]/(xy)$. A $n$-pointed curve is a $(n+1)$-tuple $(X,p_{1},\dots,p_{n})$, with $X$ a curve and $p_{1},\dots,p_{n}$ distinct smooth points of $X$.
If $X$ is a $n$-pointed nodal curve, its dual graph $\Gamma(X)$ is a graph with a vertex $v$ for each irreducible component $X^{v}_{\overline{K}}$ of a base change of $X$ to an algebraic closure $\overline{K}$, and an edge between two vertices for each node between the corresponding irreducible components. The dual graph $\Gamma(X)$ is weighted with $g(v)=g(X^{v}_{\overline{K}})$, the geometric genus of the irreducible components, and has $n$ labelled flags corresponding to the marked points.
@@
The definition is well posed, as the dual graph of a curve depends only on its topological type.
@@ 3.2. Fine compactified (universal) Jacobians
Again, the notation concerning Jacobians is taken from [PT23]. We recall the main concepts we will be using in this section.
@@
The algebraic space $\textup{Pic}^{d}(\mathcal{X}/S)$ parametrising line bundles of relative degree $d$ embeds in the algebraic space $\textup{Simp}^{d}(\mathcal{X}/S)$ parametrising rank $1$ torsion-free sheaves of degree $d$ on the family of nodal curves $\mathcal{X}\to S$, by [AK80] and [Est97].
While being locally of finite type over $S$ and satisfying the existence part of the valuative criterion for properness, in general $\textup{Simp}^{d}(\mathcal{X}/S)$ fails to be separated and of finite type.
@@
The analogous notion is defined for families of nodal curves and in particular a degree $d$ fine compactified universal Jacobian of type $(g,n)$ is an open proper substack $\overline{\mathcal{J}}_{g,n}\subset\textup{Simp}^{d}(\overline{\mathcal{C}}_{g,n}/\overline{\mathcal{M}}_{g,n})$. We denote its fibre over a curve $X$ by $\overline{\mathcal{J}}_{g,n}(X)$, dropping the subscript when it is clear from the context.
@@ 3.3. Correspondence with universal stability conditions
In this section, we combine the classification result for fine compactified universal Jacobians in [PT23, Corollary 7.13] and the correspondence between $V$-stabilities and fine compactified Jacobians in [Viv23, Theorem 1.20, Theorem 2.33], to show that unversal $c$-stability conditions bijectively correspond to fine compactified universal Jacobians. Thus, we obtain an explicit combinatorial description, thanks to which we will be able to provide examples of fine compactified universal Jacobians, whose fibres are not all classically polarised.
Let $\overline{\mathcal{J}}_{g,n}$ be a fine compactified universal Jacobian, and let $[\mathcal{F}]\in\overline{\mathcal{J}}_{g,n}$ be a family of sheaves. By [PT23, Remark 7.10], if $X$ and $X^{\prime}$ are two curves with $\Gamma(X)=\Gamma(X^{\prime})$, then $\underline{\textup{deg}}(\mathcal{F}_{|_{X}})=\underline{\textup{deg}}(\mathcal{F}_{|_{X^{\prime}})}$. In particular, for each $(e,h,A)\in\mathcal{D}_{g,n}$, the degree of $\mathcal{F}$ on the first component $C(e,h,A)^{v_{1}}$ of each vine curve of type $(e,h,A)$ is a well-defined integer, which we denote by $\textup{deg}_{C(e,h,A)^{v_{1}}}(\mathcal{F})$. Thus, we can give the following definition.
@@
More specifically, the bijection provided by [Viv23, Theorem 1.20], together with the compatibility with all the morphisms of $G_{g,n}$ induced by étale specialisations in ¯$\overline{\mathcal{M}}_{g,n}$ guaranteed by the above reasoning, maps $\mathfrak{m}(\overline{\mathcal{J}}_{g,n})$ to the associated assignment ([PT23, Definition 7.1]) $\sigma_{\overline{\mathcal{J}}_{g,n}}$ of ${\overline{\mathcal{J}}_{g,n}}$.
@@
Let $\mathfrak{m}$ be a degree $d$ universal stability condition of type $(g,n)$. We define the fine compactified universal Jacobian associated to $\mathfrak{m}$ as
@@
By Lemma 2.9, $\mathfrak{m}$ is the inverse limit of a family $\{\mathfrak{m}^{\Gamma}\}_{\Gamma\in G_{g,n}}$. For each $\Gamma\in G_{g,n}$ and for each curve $X\in\overline{\mathcal{M}}_{g,n}$ such that $\Gamma(X)=\Gamma$, the fibre $\overline{\mathcal{J}}_{g,n}(\mathfrak{m})_{|_{X}}$ coincides with the $V$-open subset $\overline{J}_{X}(\mathfrak{m}^{\Gamma})$ of [Viv23, Lemma-Definition 2.30], by Proposition 2.12, and if $\sigma(\mathfrak{m}^{\Gamma})$ is the stability condition (according to [PT23, Definition 4.3]) associated to $\mathfrak{m}^{\Gamma}$ by [Viv23, Theorem 1.20], Theorem 2.33 loc.cit. states that $\overline{J}_{X}(\mathfrak{m}^{\Gamma})=\overline{J}_{\sigma(\mathfrak{m}^{\Gamma})}$ ([PT23, Definition 6.2]). Finally, since $\mathfrak{m}_{(e,h,A)}^{\Gamma}=\mathfrak{m}_{(e,h,A)}^{\Gamma^{\prime}}$ for each $(e,h,A)\in\mathcal{D}(\Gamma)\cap\mathcal{D}(\Gamma^{\prime})$, the family $\{\sigma(\mathfrak{m}^{\Gamma})\}_{\Gamma\in G_{g,n}}$ is a family of degree $d$ stability conditions on $\overline{\mathcal{C}}_{g,n}/\overline{\mathcal{M}}_{g,n}$ ([PT23, Definition 4.7]). Therefore, by Theorem 6.3 in loc.cit., we conclude that $\overline{\mathcal{J}}_{g,n}(\mathfrak{m})$ is, indeed, a degree $d$ fine compactified universal Jacobian of type $(g,n)$. ∎
@@
Let $\mathfrak{m}$ be a degree $d$ universal stability condition of type $(g,n)$. We say that the fine compactified universal Jacobian $\overline{\mathcal{J}}_{g,n}(\mathfrak{m})$ associated to $\mathfrak{m}$ is classical if there exists a classical universal stability condition $\phi$ such that $\mathfrak{m}=\mathfrak{m}(\phi)$.
Let $[X]\in\overline{\mathcal{M}}_{g,n}$. We say that the fibre $\overline{\mathcal{J}}_{X}(\mathfrak{m}^{\Gamma(X)})$ of $\overline{\mathcal{J}}_{g,n}(\mathfrak{m})$ over $[X]$ is classical if there exists a classical stability condition $\phi^{\Gamma(X)}$ on the dual graph $\Gamma(X)$ of $X$ such that $\mathfrak{m}^{\Gamma(X)}=\mathfrak{m}_{|_{\mathcal{D}(\Gamma(X)}}=\mathfrak{m}^{\Gamma(X)}(\phi^{\Gamma(X)})$.
@@
of Theorem A.
Since both arrows are well defined by the above discussion, we need only to show they are indeed mutual inverses. Let $\overline{\mathcal{J}}_{g,n}$ be a degree $d$ fine compactified universal Jacobian. By Remark 3.3, the stability condition $\mathfrak{m}(\overline{\mathcal{J}}_{g,n})$ is mapped to the associated assignment of $\overline{\mathcal{J}}_{g,n}$, by the fibrewise application of [Viv23, Theorem 1.20]. Now, the proof of Lemma-Definition 3.5, shows that the fine compactified universal Jacobian associated to $\mathfrak{m}(\overline{\mathcal{J}}_{g,n})$, is exactly the moduli space of $\sigma(\mathfrak{m})$-stable sheaves (according to [PT23, Definition 6.2]). Hence Corollary 7.11 in loc.cit. proves that $\overline{\mathcal{J}}_{g,n}(\mathfrak{m}(\overline{\mathcal{J}}_{g,n}))=\overline{\mathcal{J}}_{g,n}$. Analogously, by the second part of the same theorem, we see that $\mathfrak{m}(\overline{\mathcal{J}}_{g,n}(\mathfrak{m}))=\mathfrak{m}$, thus concluding our proof.
Such a description allows us to show that, for each hyperbolic pair $(g,n)$, there exists a finite number of degree $d$ fine compactified universal Jacobians of type $(g,n)$, up to the translation action of the universal, relative degree 00 Picard group $\textup{PicRel}^{0}_{g,n}(\mathbb{Z})$ ([KP19, Definition 2.1]). If $g\geq 2$ and $n\geq 1$, the generators of $\textup{PicRel}^{0}_{g,n}(\mathbb{Z})$ are the line bundles $\mathcal{O}(\Sigma_{1}-\Sigma_{j})$, for $j=2,\dots,n$ and $\mathcal{O}((2g-2)\Sigma_{1})\otimes\omega_{\pi}^{\otimes-1}$, where $\Sigma_{j}$ denotes the $j$-th section of the universal curve $\pi:\overline{\mathcal{C}}_{g,n}\to\overline{\mathcal{M}}_{g,n}$, and $\omega_{\pi}$ is the relative dualising sheaf.
@@
The degree $d$ universal stability conditions $\mathfrak{m}$ on of type $(g,n)$, with $g\geq 2$ and $n\geq 1$, such that $\mathfrak{m}_{(2,0,\{1\})}\in\{0,\dots,2g-3\}$ and $\mathfrak{m}_{(2,0,\{i\})}=0$ for each $i\neq 1$ form a full set of representatives for the action of $\textup{PicRel}^{0}_{g,n}(\mathbb{Z})$ on the set of degree $d$ universal stability conditions $\mathfrak{m}$ of type $(g,n)$.
@@
Let $\mathfrak{m}$ be a universal stability condition. First, we want to show that $\mathfrak{m}$ is identified by a set of differences with its elements of the form $\mathfrak{m}_{(2,0,\{i\})}$, for $i=1,\dots,n$. The statement is clearly true for $\mathfrak{m}_{(2,0,A)}$, where $A\neq\emptyset$. In fact, the iterate application of property (2.2), shows that $\mathfrak{m}_{(2,0,A)}$ is uniquely determined by
@@
of Theorem B, 2..
Clearly, given an assignment $\{\mathfrak{n}_{A}\}_{\emptyset\neq A\subset[n]}$, there exist only a finite amount of degree $d$ universal stability conditions $\mathfrak{m}$ of type $(g,n)$ such that $\mathfrak{m}_{(2,0,A)}=\mathfrak{n}_{A}$ for each $\emptyset\neq A\subset[n]$. This proves the statement for each pair $(g,n)$, with $g\geq 2$ and $n\geq 1$.
In the case of genus $g=0$ the statement is trivial, while for $g=1$ it follows from [PT22, Theorem 6.5].
By [PT23, Theorem 9.6], every universal stability condition of type $(g,0)$ is induced by a classical universal $\phi$-stability condition. Since, by [KP19, Remark 5.11], the classical stability polytopes are finite up to translation, we conclude. ∎
@@
A further consequence of Lemma 3.7 and part two of Theorem B is that degree $d$ fine compactified universal Jacobians of type $(g,n)$ are finite up to isomorphisms that preserve to the forgetful morphism to ¯$\overline{\mathcal{M}}_{g,n}$. In fact, [KP19, Lemma 6.13], with $L$ of degree 00 and $t=0$ if $d\neq 0$, holds true with the same proof.

### v2

@@ 2.2. Universal and c-stability conditions
By Proposition 2.17, $\mathfrak{m}_{(e,h,A)}^{\Gamma}=\mathfrak{m}_{(e,h,A)}^{\Gamma}(\phi^{\Gamma})$ if and only if
(2.7) $\Bigl{\lceil}\frac{(e+2h-2)(d+1-g)+(2g-e-2h)\phi^{\Gamma}_{A}-(e+2h-2)\phi^{\Gamma}_{A^{\mathsf{c}}}}{2g-2}\Bigr{\rceil}+h-1=\mathfrak{m}_{(e,h,A)}^{\Gamma},$
while $\mathfrak{m}^{\Gamma}_{(e,g+1-e-h,A^{\mathsf{c}})}=\mathfrak{m}^{\Gamma}_{(e,g+1-e-h,A^{\mathsf{c}})}(\phi^{\Gamma})$ if and only if
(2.8) $\Bigl{\lceil}\frac{(-e+2g-2h)(d+1-g)+(e+2h-2)\phi^{\Gamma}_{A^{\mathsf{c}}}-(2g-e-2h)\phi^{\Gamma}_{A}}{2g-2}\Bigr{\rceil}+g-e-h=\mathfrak{m}^{\Gamma}_{(e,g+1-e-h,A^{\mathsf{c}})}.$
@@
Clearly, if we choose a family $\{\phi^{\Gamma}\}_{\Gamma\in G_{g,n}}$ defining a universal numerical polarisation, the same proof applied to all $(e,h,A)\in\mathcal{D}_{g,n}$ shows that we can check if a universal stability condition is classical on half the vertices of the vine graphs.
@@ 3.1. Curves and dual graphs
We recall the curve-theoretic notation we will use throughout the paper. We refer to [ACGH13] and [PT] for a detailed exposition.
@@
A curve over an extension $K$ of $k$ is a $\textup{Spec}(K)$-scheme $X$ proper over $K$, geometrically connected and of pure dimension $1$. A curve $X$ is nodal if it is geometrically reduced and, when passing to an algebraic closure $\overline{K}$, its local ring at every singular point is isomorphic to $\overline{K}[[x,y]]/(xy)$. A $n$-pointed curve over $K$ is a $(n+1)$-tuple $(X,p_{1},\dots,p_{n})$, with $X$ a curve and $p_{1},\dots,p_{n}$ distinct smooth $K$-points of $X$.
If $X$ is a $n$-pointed nodal curve, its dual graph $\Gamma(X)$ is a graph with a vertex $v$ for each irreducible component $X^{v}_{\overline{K}}$ of a base change of $X$ to an algebraic closure $\overline{K}$, and an edge between two vertices for each node, defined over $\overline{K}$, between the corresponding irreducible components. The dual graph $\Gamma(X)$ is weighted with $g(v)=g(X^{v}_{\overline{K}})$, the geometric genus of the irreducible components, and has $n$ labelled flags corresponding to the marked points.
@@ 3.2. Fine compactified (universal) Jacobians
Again, the notation concerning Jacobians is taken from [PT]. We recall the main concepts we will be using in this section.
@@
The algebraic space $\textup{Pic}^{d}(\mathcal{X}/S)$ parametrising line bundles of relative degree $d$ embeds in the algebraic space $\textup{Simp}^{d}(\mathcal{X}/S)$ parametrising rank $1$ torsion-free simple sheaves of degree $d$ on the family of nodal curves $\mathcal{X}\to S$, by [AK80] and [Est01].
While being locally of finite type over $S$ and satisfying the existence part of the valuative criterion for properness, as proven in [Est01], in general $\textup{Simp}^{d}(\mathcal{X}/S)$ fails to be separated and of finite type.
@@
The analogous notion is defined for families of nodal curves and in particular a degree $d$ fine compactified universal Jacobian of type $(g,n)$ is an open proper substack $\overline{\mathcal{J}}_{g,n}\subset\textup{Simp}^{d}(\overline{\mathcal{C}}_{g,n}/\overline{\mathcal{M}}_{g,n})$ (which is automatically connected). We denote its fibre over a curve $X$ by $\overline{\mathcal{J}}_{g,n}(X)$, dropping the subscript when it is clear from the context.
@@ 3.3. Correspondence with universal stability conditions
In this section, we combine the classification result for fine compactified universal Jacobians in [PT, Corollary 7.13] and the correspondence between $V$-stabilities and fine compactified Jacobians in [Viv, Theorem 1.20, Theorem 2.33], to show that universal stability conditions bijectively correspond to fine compactified universal Jacobians. Thus, we obtain an explicit combinatorial description, thanks to which we will be able to provide examples of fine compactified universal Jacobians that are not semi-classical, i.e. whose fibres are not all classical fine compactified Jacobians.
Let $\overline{\mathcal{J}}_{g,n}$ be a fine compactified universal Jacobian, and let $[\mathcal{F}]\in\overline{\mathcal{J}}_{g,n}$ be a family of sheaves. By [PT, Remark 7.10], if $X$ and $X^{\prime}$ are two curves with $\Gamma(X)=\Gamma(X^{\prime})$, then $\underline{\textup{deg}}(\mathcal{F}_{|_{X}})=\underline{\textup{deg}}(\mathcal{F}_{|_{X^{\prime}})}$. In particular, for each $(e,h,A)\in\mathcal{D}_{g,n}$, the degree of $\mathcal{F}$ on the first component $C(e,h,A)^{v_{1}}$ of each vine curve of type $(e,h,A)$ is a well-defined integer, which we denote by $\textup{deg}_{C(e,h,A)^{v_{1}}}(\mathcal{F})$. Thus, we can give the following definition.
@@
More specifically, the bijection provided by [Viv, Theorem 1.20], together with the compatibility with all the morphisms of $G_{g,n}$ induced by étale specialisations in ¯$\overline{\mathcal{M}}_{g,n}$ guaranteed by the above reasoning, maps $\mathfrak{m}(\overline{\mathcal{J}}_{g,n})$ to the associated assignment ([PT, Definition 7.1]) $\sigma_{\overline{\mathcal{J}}_{g,n}}$ of ${\overline{\mathcal{J}}_{g,n}}$.
@@
Let $\mathfrak{m}$ be a degree $d$ universal stability condition of type $(g,n)$. The fine compactified universal Jacobian associated to $\mathfrak{m}$ as
@@
By Lemma 2.9, $\mathfrak{m}$ is the inverse limit of a family $\{\mathfrak{m}^{\Gamma}\}_{\Gamma\in G_{g,n}}$. For each $\Gamma\in G_{g,n}$ and for each curve $X\in\overline{\mathcal{M}}_{g,n}$ such that $\Gamma(X)=\Gamma$, the fibre $\overline{\mathcal{J}}_{g,n}(\mathfrak{m})_{|_{X}}$ coincides with the $V$-open subset $\overline{J}_{X}(\mathfrak{m}^{\Gamma})$ of [Viv, Lemma-Definition 2.30], by Proposition 2.12. Moreover, if $\sigma(\mathfrak{m}^{\Gamma})$ is the stability condition (according to [PT, Definition 4.3]) associated to $\mathfrak{m}^{\Gamma}$ by [Viv, Theorem 1.20], [Viv, Theorem 2.33] states that $\overline{J}_{X}(\mathfrak{m}^{\Gamma})=\overline{J}_{\sigma(\mathfrak{m}^{\Gamma})}$ ([PT, Definition 6.2]).
Finally, since $\mathfrak{m}_{(e,h,A)}^{\Gamma}=\mathfrak{m}_{(e,h,A)}^{\Gamma^{\prime}}$ for each $(e,h,A)\in\mathcal{D}(\Gamma)\cap\mathcal{D}(\Gamma^{\prime})$, the family $\{\sigma(\mathfrak{m}^{\Gamma})\}_{\Gamma\in G_{g,n}}$ is a family of degree $d$ stability conditions on $\overline{\mathcal{C}}_{g,n}/\overline{\mathcal{M}}_{g,n}$ ([PT, Definition 4.7]). Therefore, by [PT, Theorem 6.3], we conclude that $\overline{\mathcal{J}}_{g,n}(\mathfrak{m})$ is a degree $d$ fine compactified universal Jacobian of type $(g,n)$. ∎
@@
Let $\mathfrak{m}$ be a degree $d$ universal stability condition of type $(g,n)$. We say that the fine compactified universal Jacobian $\overline{\mathcal{J}}_{g,n}(\mathfrak{m})$ associated to $\mathfrak{m}$ is classical (resp. semi-classical) if $\mathfrak{m}$ is classical (resp. semi-classical).
Let $[X]\in\overline{\mathcal{M}}_{g,n}$. We say that the fibre $\overline{\mathcal{J}}_{X}(\mathfrak{m}^{\Gamma(X)})$ of $\overline{\mathcal{J}}_{g,n}(\mathfrak{m})$ over $[X]$ is classical if the stability condition $\mathfrak{m}^{\Gamma(X)}=\mathfrak{m}_{|_{\mathcal{D}(\Gamma(X))}}$ on the dual graph $\Gamma(X)$ is classical.
@@
Proof of Theorem A.
Since both arrows are well defined by the above discussion, we need only to show they are indeed mutual inverses. Let $\overline{\mathcal{J}}_{g,n}$ be a degree $d$ fine compactified universal Jacobian. By Remark 3.3, the bijections given by [Viv, Theorem 1.20] map the stability condition $\mathfrak{m}(\overline{\mathcal{J}}_{g,n})$ to the associated assignment of $\overline{\mathcal{J}}_{g,n}$. Now, the proof of Lemma-Definition 3.5, shows that the fine compactified universal Jacobian associated to $\mathfrak{m}(\overline{\mathcal{J}}_{g,n})$, is exactly the moduli space of $\sigma(\mathfrak{m})$-stable sheaves (according to [PT, Definition 6.2]). Hence [PT, Corollary 7.19] proves that $\overline{\mathcal{J}}_{g,n}(\mathfrak{m}(\overline{\mathcal{J}}_{g,n}))=\overline{\mathcal{J}}_{g,n}$. Analogously, by the second part of the same corollary, we see that $\mathfrak{m}(\overline{\mathcal{J}}_{g,n}(\mathfrak{m}))=\mathfrak{m}$, thus concluding our proof.
Such a description allows us to show that, for each hyperbolic pair $(g,n)$, there exists a finite number of degree $d$ fine compactified universal Jacobians of type $(g,n)$, up to the translation action of the universal, relative degree 00 Picard group $\textup{PicRel}^{0}_{g,n}(\mathbb{Z})$ ([KP19, Definition 2.1]). If $g\geq 2$ and $n\geq 1$, the generators of $\textup{PicRel}^{0}_{g,n}(\mathbb{Z})$ are:
  • the boundary divisors $\mathcal{O}(C(1,h,A)^{v_{1}})$, for any $(1,h,A)\in\mathcal{D}_{g,n}$;
  • the line bundles $\mathcal{O}(\Sigma_{1}-\Sigma_{j})$, for $j=2,\dots,n$;
  • $\mathcal{O}((2g-2)\Sigma_{1})\otimes\omega_{\pi}^{\otimes-1}$.
Here $\Sigma_{j}$ denotes the $j$-th section of the universal curve $\pi:\overline{\mathcal{C}}_{g,n}\to\overline{\mathcal{M}}_{g,n}$, and $\omega_{\pi}$ is the relative dualising sheaf.
@@
$\mathcal{O}(C(1,h^{\prime},A^{\prime})^{v_{1}})\cdot\mathfrak{m}_{(e,h,A)}$ =$=\begin{cases*}\mathfrak{m}_{(e,h,A)}-1,\text{ if }(e,h,A)=(1,h^{\prime},A^{\prime})\\ \mathfrak{m}_{(e,h,A)}+1,\text{ if }(e,h,A)=(1,g-h^{\prime},[n]\setminus A^{\prime})\\ \mathfrak{m}_{(e,h,A)},\text{ otherwise}\end{cases*};$
@@
The degree $d$ universal stability conditions $\mathfrak{m}$ on of type $(g,n)$, with $g\geq 2$ and $n\geq 1$, such that:
  • $\mathfrak{m}_{(1,h,A)}=\begin{cases*}0,\textup{ if }1\in A\\ d,\textup{ if }1\notin A\end{cases*}$, for each $(1,h,A)\in\mathcal{D}_{g,n}$;
  • $\mathfrak{m}_{(2,0,\{1\})}\in\{0,\dots,2g-3\}$;
  • $\mathfrak{m}_{(2,0,\{i\})}=0$ for each $i\neq 1$;
form a full set of representatives for the action of $\textup{PicRel}^{0}_{g,n}(\mathbb{Z})$ on the set of degree $d$ universal stability conditions $\mathfrak{m}$ of type $(g,n)$.
@@
Let $\mathfrak{m}$ be a universal stability condition. Applying $\otimes\mathcal{O}(C(1,h,A)^{v_{1}})^{\otimes\mathfrak{m}_{(1,h,A)}}$, for each $(1,h,A)\in\mathcal{D}_{g,n}$ such that $1\in A$, we obtain a translation equivalent stability condition that satisfies the first requirement.
Now, let us consider the latter two conditions. First, we want to show that $\mathfrak{m}$ is identified by a set of differences with its elements of the form $\mathfrak{m}_{(2,0,\{i\})}$, for $i=1,\dots,n$. The statement is clearly true for $\mathfrak{m}_{(2,0,A)}$, where $A\neq\emptyset$. In fact, the iterate application of property (2.2), shows that $\mathfrak{m}_{(2,0,A)}$ is uniquely determined by
@@
Proof of Theorem B, 2..
Clearly, given an assignment $\{\mathfrak{n}_{A}\}_{\emptyset\neq A\subset[n]}$, there exist only a finite amount of degree $d$ universal stability conditions $\mathfrak{m}$ of type $(g,n)$ such that $\mathfrak{m}_{(2,0,A)}=\mathfrak{n}_{A}$ for each $\emptyset\neq A\subset[n]$, as they need to satisfy (2.2). Since the assignments on the triples $(1,h,A)$ are fixed, this proves the statement for each pair $(g,n)$, with $g\geq 2$ and $n\geq 1$.
In the case of genus $g=0$ the statement is trivial, while for $g=1$ it follows from [PT23, Theorem 6.5].
By [PT, Theorem 9.8], there is at most one degree $d$ universal stability condition of type $(g,0)$, for any genus $g$. Thus, we conclude. ∎
@@
A further consequence of Lemma 3.7 and part two of Theorem B is that degree $d$ fine compactified universal Jacobians of type $(g,n)$ are finite up to isomorphisms that commute with the forgetful morphism to ¯$\overline{\mathcal{M}}_{g,n}$. In fact, [KP19, Lemma 6.13], with $L$ of degree 00 and $t=0$ if $d\neq 0$, holds true with the same proof.